( * ) If A,B,C are acture angles such that sin (B+C-A) = cos(C+A-B) = tan(A+B-C) = 1 then (A,B,C)=

( * )
If A,B,C are acture angles such that sin (B+C-A) = cos(C+A-B) = tan(A+B-C) = 1 then (A,B,C)=

Grade:11

1 Answers

Nishant Vora IIT Patna
9 years ago
Hello student, Please find the answer to your question below

sin (B+C-A) =1 => B+C-A = 90 ….….(1)
cos(C+A-B) = 1 => C+A-B= 0….…..(2)
tan(A+B-C) = 1 => A+B-C = 45….….(3)

Now there are three equations and three variable i.e A,B,C ..Easily solvabe